Fermi National Laboratory


Accelerator Update for March 30, 2005 - April 1, 2005

Wednesday March 30
The day shift began with Operations monitoring store 4063 and the Recycler stash, with the Antiproton Source (Pbar) stacking, and with MiniBooNE and the Meson MCenter experiments taking beam.

Operations terminated store 4063 at 9:04 AM. Operators began preparing the Tevatron (TeV) and the Collider Detector at Fermilab (CDF) and D-Zero (D0) collision halls for access.

Cryo system technicians began work on the TeV sector E3 wet engine at 9:13 AM.

LCW system techs began work at CUB.

Operations established beam at 10:16 AM to the Meson experiment MTest.

Operations began setting up NuMI at 10:44 AM for beam to scan their target.

At 11:07 AM, a NuMI power supply (E:HV100) tripped off. Operators contacted experts who had the supply back on by 12:04 PM.

Operations put the TeV into shot setup at 1 PM.

Operations established combination store 4065 at 3:41 PM with an initial luminosity of 85.46E30. Pbar resumed stacking and MCenter and MTest resumed taking beam.

NuMI experts began conducting target scans at 4:13 PM.

At 5 PM, Operations reported the loss of the NuMI beam permit. Experts investigated.

At 5:01 PM, a Recycler power supply (CE511) tripped off. An expert investigated and began replacing one of the supplies.

At 5:18 PM, Operations first reported losing the Main Injector (MI) power supply beam permit, then they reported that Fermilab had suffered a power glitch that aborted store 4065. Linac, Booster, and MI are all down due to many device trips.

At 7:03 PM, cryo techs at CHL reported losing one helium compressor. There are still two compressors operating so the system is stable.

Operations reported establishing a MI ramp at 7:34 PM.

Pbar resumed stacking at 7:47 PM.

At 7:53 PM, a Switchyard (SY) power supply (S:HP3US) developed an LCW leak. There will be no beam to SY until this gets fixed.

Operations established beam to NuMI at 8:25 PM.

TeV experts began a study period at 8:52 PM.

The High Voltage expert reported at 9:05 PM that the power glitch was from off site.

NuMI experts completed their target scans at 9:57 PM.

Thursday March 31
The midnight shift began with TeV expert conducting studies, with SY off due to a HP3US LCW leak, with Pbar stacking, and with MiniBooNE taking beam.

Operations loaded a 36x0 study-store into the TeV at 12:10 AM.

Operations reported that techs had completed HP3US repairs at 12:25 AM.

At 12:27 AM, Operations reported losing the beam permit for the P150 transfer line for the first of three times in an hour.

At 1:35 AM, a MI power supply (V701), located in the P1 transfer line, tripped off and wouldn't reset. Operators called back in the power supply techs. They had the supply repaired by 3:12 AM. Pbar resumed stacking and TeV resumed their studies.

Operators and Safety experts prepared MI-65 for a NuMI access at 8:06 AM.

Operators prepared the Booster for a short access at 9:40 AM to work on a blower for Booster RF 13.

Operations loaded at 4x0 study-store for a wet squeeze at 10:39 PM.

At 11:32 AM, Operations reported a TeV TECAR problem. Experts investigated the problem.

Operations reported at 12:08 PM that the TeV TECAR experts had completed their work and requested a dry squeeze followed by a wet squeeze.

Operations loaded a 4x0 study-store into the TeV at 1:02 PM.

Operations put the TeV into shot setup at 1:37 PM.

At 2:06 PM, the MiniBooNE beam permit dropped and Operations received a smoke alarm for MI-10. The problem was with a magnet (HV860). It was also noted that the MI LCW system was losing water. Operators began preparing MI for an access at 2:20 PM.

Operators prepared the Booster for an access at 2:28 PM.

At 2:50 PM, experts at MI-10 reported finding a leak in a LCW line for one of the HV860 magnets. They valved out the magnet. The magnet had overheated at a power flag and scorched the magnet coil. Expert continued to investigate.

scorched magnet coil

At 4:52 PM, Operations received a report of a problem in the Booster East Gallery. Operators and experts investigated and later discovered a quad power supply (B:QDB1) had failed. They replaced the supply.

MI resumed accelerating beam at 6 PM.

Operations put the TeV into shot setup at 7:09 PM. The antiprotons will only come from Pbar for this shot.

Operations established store 4072 at 10:24 PM with an initial luminosity of 65.47E30. Pbar resumed stacking.

Friday April 1
The midnight shift began with Operations monitoring store 4072, with Pbar stacking, with MiniBooNE off due to magnet trouble, and with MCenter taking beam.

At 5:11 AM, Operations reported a loss in luminosity. Operators investigated and found a damper problem.

The Plan for Friday and the Weekend
The plan for today is to determine if HV860 can be repaired in place, or if the magnet must be replaced. The Run Coordinator has developed several scenarios depending on what might occur, which might be anywhere from turning back on in a few hours after repairs, to replacing the magnet and turning on late Sunday or Monday.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 4/4/2005   email Fermilab